# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v2

@@ Abstract
Here, $U(t)$ is the time evolution operator. Thus we have a correspondence between multi-tone driven localized phases and parameterized families of QCA, $V({\vec{\theta}})$. That is, homotopy classes of continuous maps from the $n$-dimensional torus, ${\mathbb{T}}^{n}:=[{\mathbb{R}}/(2\pi{\mathbb{Z}})]^{n}$, to $\mathrm{QCA}_{d}$, which we denote $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$. Ignoring weak invariants [100, 101] corresponding to low dimensional cross sections of ${\mathbb{T}}^{n}$, we may instead consider the $n$th homotopy group $\pi_{n}(\mathrm{QCA}_{d})$.
Once more, there is an induced homomorphism $e^{*}_{0}:\pi_{n}(\mathrm{QCA}_{d})\to\pi_{n}(\mathrm{inv}_{d})$, and ALT phases are identified as the nontrivial elements of $\ker e^{*}_{0}$. As in the case of static systems, the parameterization by QCA is not unique, so that the classification of ALT phases is properly a quotient of $\ker e^{*}_{0}$. Further details regarding driven MBL phases are provided throughout Appendix A.
Classification.—The classification of static ALT phases in $d$ dimensions is determined by the classification of QCA. That is, by $\pi_{0}(\mathrm{QCA}_{d})$. There has been great recent progress in classifying QCA in spin systems—a complete classification now exists in one [58] and two [61] dimensions, and several nontrivial phases are known in three dimensions [56, 55, 57]. The classification of ALT phases inherits this progress.
For Floquet systems, there is another correspondence known between localized phases and QCA. Reference [23] showed that a periodically driven localized phase can be classified by its action on the edge when placed on open boundary conditions. This action is described by a QCA of one dimension lower, $\pi_{0}(\mathrm{QCA}_{d-1})$. On the other hand, both our construction and an equivalent formulation of unitary loops [49] show that Floquet localized phases should be classified by $\pi_{1}(\mathrm{QCA}_{d})$. We deduce that these groups must be isomorphic,
@@
The relationship Eq. (4) is reminiscent of a widely held conjecture that invertible states form an $\Omega$-spectrum [96, 95, 102, 103, 104]. That is, there is a homotopy equivalence [105] $\mathrm{inv}_{d-1}\simeq\Omega\mathrm{inv}_{d}$, where $\Omega X$ is the space of loops that begin and end at a specified base point in the space $X$. In particular $\pi_{n-1}(\mathrm{inv}_{d-1})\cong\pi_{n}(\mathrm{inv}_{d})$ for all $n$. We make the conjecture that QCA also forms an $\Omega$-spectrum,
@@
The base point for the loops in $\Omega\mathrm{QCA}_{d}$ is the identity.
@@
Equation (5) is interpreted through a picture of pumping [95, 104, 103]. Taking a loop $\gamma\in\Omega\mathrm{QCA}_{d}$ and restricting it to open boundary conditions, the action in the bulk must still be trivial at the end of the loop, due to the finite range of the QCA. However, on the $(d-1)$-dimensional edge, there may be a nontrivial action by some $v\in\mathrm{QCA}_{d-1}$ [23]. We say that $\gamma$ pumps $v$ across the system, and Eq. (5) hypothesizes that the homotopy structure of the loop space $\Omega\mathrm{QCA}_{d}$ is completely characterized by the pumped QCA $v$.
More formally, we argue that the swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ is a homotopy equivalence. Given some $v\in\mathrm{QCA}_{d-1}$, the swindle $S(v,t)$ is defined through a finite depth quantum circuit [95, 106, 60, 59] (Fig. 2). Let $\mathrm{swap}_{jk}(t)$ be a path from the identity to the swap gate in the unitary group on sites $j$ and $k$. We stack these gates together in a codimension 1 plane,
$\mathrm{SWAP}_{x,x+1}(t)=\bigotimes_{\vec{y}\in{\mathbb{Z}}^{d-1}}\mathrm{swap}_{(x,\vec{y}),(x+1,\vec{y})}(t).$ (6)
Then the path of QCA on a two-layer system
@@
interpolates between the identity and $v_{x}v_{x+1}^{\dagger}$, where $v_{x}$ acts as $v$ on the $(d-1)$-dimensional surface $\{x\}\times{\mathbb{Z}}^{d-1}$. Finally, the swindle map is
$S(v,t)=\left(\prod_{x\in 2{\mathbb{Z}}+1}u_{x,x+1}(v,t)\right)\left(\prod_{x\in 2{\mathbb{Z}}}u_{x,x+1}(v,t)\right),$ (8)
which is continuous in $t$ and $v$ and equal to the identity for $(t,v)\in\{0,1\}\times\mathrm{QCA}_{d-1}$ (Fig. 2).
Assuming that any loop of QCA can be restricted to open boundary conditions such that it remains a QCA, we show that $S$ is a homotopy equivalence in Appendix B. Formal details are deferred to Ref. [107].
Iteratively applying Eq. (5), we reduce the classification of driven localized phases in any dimension to either the classification of static QCA, or a driven classification of zero-dimensional QCA,
@@
The space $\mathrm{QCA}_{0}$ is a stabilization of the projective unitary group by stacking ancillae. We find its homotopy groups in Appendix C. Using the known classification of one- and two-dimensional QCA, we have all the homotopy groups $\pi_{n}(\mathrm{QCA}_{d})$ for spin systems with $d-n\leq 2$. These are listed in Tab. 1, and the corresponding classification of localized phases (excluding permutation QCA) is listed in Tab. 2. Comparing to the classification of invertible states shows that all the nontrivial phases in the tables are ALT phases.
$q$ $3$ $2$ $1$ 00 $-1$ $-2$ $-3$ $-4$ $-5$
$h^{q}$ 00 $\log{\mathbb{Q}}^{\times}$ 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00 ${\mathbb{Q}}$ 00 ${\mathbb{Q}}$
$h_{\mathrm{inv}}^{q}$ 00 ${\mathbb{Z}}$ 00 00 00 ${\mathbb{Z}}$ 00 00 00
@@
Table 1: The homotopy groups $\pi_{n}(\mathrm{QCA}_{d})=:h^{q}$ are a function of $q=d-n$, as is $\pi_{n}(\mathrm{inv}_{d})=:h_{\mathrm{inv}}^{q}$. For spin systems $h^{3}$ is conjectured to be the Witt group of modular tensor categories [56, 108, 55, 109, 57]; $h^{2}$ is trivial [61]; $h^{1}$ is isomorphic to $\log{\mathbb{Q}}^{\times}$ [58]; and $h^{q}$ for $q\leq 0$ is trivial for even $q$, ${\mathbb{Q}}/{\mathbb{Z}}$ for $q=-1$, and ${\mathbb{Q}}$ for odd $q\leq-3$ (Appendix C). In all groups in the table, the group operation is addition.
$d\;\;$ $\;\;n$ 00 $1$ $2$ $3$ $4$ $5$
00 00 00 00 ${\mathbb{Q}}$ 00 ${\mathbb{Q}}$
$1$ 00 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00 ${\mathbb{Q}}$ 00
$2$ 00 $\log{\mathbb{Q}}^{\times}$ 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00 ${\mathbb{Q}}$
$3$ 00 $\log{\mathbb{Q}}^{\times}$ 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00
@@
Table 2: The classification of strong invariants in $n$-tone driven $d$-dimensional localized phases of spins with short-range entanglement. The classifying group differs from $\pi_{n}(\mathrm{QCA}_{d})$ for $(d,n)\in\{(0,1),(1,0)\}$ due to the non-uniqueness of the parameterization by QCA. All nontrivial elements of this table correspond to ALT phases.

### v3

@@ II.3 Driven phases
Here, $U(t)$ is the time evolution operator, the Floquet Hamiltonian $H_{F}=\sum_{j}h_{j}\sigma^{z}_{j}+\cdots$ is trivially MBL, and the micromotion operator $V({\vec{\theta}})$ is a family of QCA parameterized by $\vec{\theta}$}. Thus, in complete analogy with the static case, we have a correspondence between multi-tone-driven localized phases and parameterized families of QCA, $V({\vec{\theta}})$, up to continuous deformations. That is, homotopy classes of continuous maps from the $n$-dimensional torus to $\mathrm{QCA}_{d}$. This reduces to the $n$th homotopy group $\pi_{n}(\mathrm{QCA}_{d})$ when ignoring weak invariants [105, 106] corresponding to low dimensional cross sections of the torus. Collapsing some of these cross-sections to a point reduces the torus to a sphere, and the $n$th homotopy group precisely corresponds to families $V({\vec{\theta}})$ parameterized by a variable on a sphere, up to continuous deformations.
Once more, there is an induced homomorphism
$e^{*}_{0}:\pi_{n}(\mathrm{QCA}_{d})\to\pi_{n}(\mathrm{inv}_{d}).$ (5)
The group $\pi_{n}(\mathrm{inv}_{d})$ can be interpreted as the classification of individual quasienergy states (states for which time evolution synchronizes with the drive). The map $e_{0}^{*}$ identifies the eigenstate order associated to individual quasienergy states, and ALT phases are again identified as the nontrivial elements of $\ker e^{*}_{0}$.
As in the case of static systems, the parameterization by QCA is not unique, so that the classification of ALT phases is properly a quotient of $\ker e^{*}_{0}$. Further details are provided in Appendix A.
III Classification
[…]
In this section, we use these known facts about the classification of QCA, and argue for new results regarding the higher homotopy groups of QCA, to compute the classification of both static and driven MBL phases in low dimensions. Namely, we propose and motivate a conjecture that QCA form an $\Omega$-spectrum, which can be viewed as a bulk-boundary correspondence between driven MBL phases and phases with one less drive in one lower dimension (Sec. III.1). This conjecture is supported by an explicit construction mapping QCAs in d1d-1 dimensions to loops of QCAs (driven ALT phases) in one higher dimension, called the swindle map (Sec. III.2). Assuming the $\Omega$-spectrum conjecture is true, we calculate the complete classification of $n$-tone-driven ALT phases in $d$-dimensions with $d-n\leq 2$ (Sec. III.3).
III.1 Loops and $\Omega$-spectra
In terms of the correspondence between MBL phases and QCA in Sec. II, driven systems can be understood as relating to loops or tori of QCA. Higher homotopy groups of QCA have received little attention in the literature (though, Ref. [49] deals with essentially the same context in a different langauge), and these will be the subject of most of our new mathematical results. In particular, we propose that QCA form a mathematical object known as an $\Omega$-spectrum [101, 100, 63, 64, 65], which in more plain language means that there is a bulk-boundary correspondence between $n$-tone driven phases in $d$ dimensions and (n1)(n-1)-tone driven phases in d1d-1 dimensions.
We motivate this proposal by recalling that, for Floquet systems $(n=1)$, there is another correspondence, distinct to that of Sec. II.3, known between localized phases and QCA. Reference [23] showed that a periodically driven localized phase can be classified by its action on the edge when placed on open boundary conditions. This action is described by a QCA of one dimension lower, $\pi_{0}(\mathrm{QCA}_{d-1})$. On the other hand, both our construction and Ref. [49] show that Floquet localized phases should be classified by $\pi_{1}(\mathrm{QCA}_{d})$. We propose that these groups are isomorphic,
@@ III.1 Loops and $\Omega$-spectra
The relationship Eq. (6) is reminiscent of a widely held conjecture that invertible states form an $\Omega$-spectrum [101, 100, 63, 64, 65]. In topology, an $\Omega$-spectrum refers to a sequence of topological spaces $(Y_{d})_{d\in{\mathbb{Z}}}$ together with (weak) homotopy equivalences (that is, continuous maps which induce isomorphisms on all homotopy groups)
$Y_{d-1}\simeq\Omega Y_{d}$ (7)
where $\Omega Y_{d}$ is the space of loops in $Y_{d}$ which begin and end at a specified base point. That is, $\Omega Y_{d}$ is the space of maps $\gamma:[0,1]\to Y_{d}$ such that $\gamma(0)=\gamma(1)=*_{d}\in Y_{d}$ is the base point of $Y_{d}$. Thus, continuous paths in $\Omega Y_{d}$ correspond to homotopies of loops, and so $\pi_{0}(Y_{d-1})\cong\pi_{0}(\Omega Y_{d})\cong\pi_{1}(Y_{d})$. In fact, a similar equation holds for all higher homotopy groups, $\pi_{n-1}(Y_{d-1})=\pi_{n}(Y_{d})$. Iterating this equation shows that the homotopy groups $\pi_{n}(Y_{d})$ only depend on q=dnq=d-n. Thus, knowing all the connected components $\pi_{0}(Y_{d})$ already provides the classification of all higher homotopy groups.
$\Omega$-spectra play a special role in topology, as they serve to define generalized cohomology theories [100, 63, 65, 64]. Given the prevalence of cohomology classifications in the study of invertible topological states, it should be plausible that generalized cohomology may find application in this problem, as indeed it does. Explicitly, the $d$th generalized cohomology group of a space $X$ with coefficients in the $\Omega$-spectrum $(Y_{d})_{d\in{\mathbb{Z}}}$ is defined to be
$h^{d}_{Y}(X)=\pi_{0}\left[X,Y_{d}\right]$ (8)
(where $\left[X,Y_{d}\right]$ is the space of continuous functions from $X$ to $Y_{d}$).
The hypothesis that invertible states form an $\Omega$-spectrum implies immediately that $\pi_{n}(\mathrm{inv}_{d})=\pi_{n-1}(\mathrm{inv}_{d-1})$. Further, it is believed that the classification of symmetry-enriched invertible states with symmetry group $G$ are completely classified by the generalized cohomology of the classifying space $BG$ [i.e. $h^{d}_{\mathrm{inv}}(BG)$[100, 63, 65, 64].
@@
The base point for the loops in $\Omega\mathrm{QCA}_{d}$ is the identity. For $d<0$, we define $\mathrm{QCA}_{d}=\Omega^{-d}\mathrm{QCA}_{0}$, making the negative degrees of this statement trivial.
We unravel the physical interpretation of Eq. (9) in terms of a bulk-boundary correspondence through a picture of pumping [100, 65, 64]. Consider a loop $\gamma\in\Omega\mathrm{QCA}_{d}$. If the path $\gamma(t)$ is differentiable, it can be realized by Hamiltonian evolution under some $H(t)$. Then we can restrict $\gamma$ to open boundary conditions by taking the partial trace of $H(t)$ outside of the specified boundaries. The path of QCA generated by this partial trace, $\tilde{\gamma}(t)$, reproduces the action of $\gamma(t)$ far from the boundary. In particular, $\tilde{\gamma}(1)$ acts as the identity on any local operator which is supported far from the boundaries. However, on the (d1)(d-1)-dimensional edge, $\tilde{\gamma}(1)$ may have a nontrivial action described by some $v\in\mathrm{QCA}_{d-1}$ [23]. We say that $\gamma$ pumps $v$ across the system. Equation (9) hypothesizes that the loop space $\Omega\mathrm{QCA}_{d}$ is characterized (up to continuous deformations) entirely by the pumped QCA $v$.
Indeed, continuous deformations of the bulk loop $\gamma$ produce continuous deformations of $v$ [107], so the topological invariants of $v$ may also be considered invariants for $\gamma$. The difficult steps needed to prove Eq. (9) are to show that:
  1. 1.
    all $v\in\mathrm{QCA}_{d-1}$ can occur as edge actions of some $\gamma\in\Omega\mathrm{QCA}_{d}$, and
  2. 2.
    every topological feature of the loop space is captured by the edge action.
III.2 The swindle map
The swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ demonstrates that 1 is true. Given some $v\in\mathrm{QCA}_{d-1}$, the swindle $S(v,t)\in\Omega\mathrm{QCA}_{d}$ pumps $v$ to the edge.
@@ III.2 The swindle map
We can construct the swindle through a finite depth quantum circuit as follows [100, 108, 60, 59] (Fig. 3). Let $\mathrm{SWAP}_{x,x+1}(t)$ be a path of finite depth circuits from the identity to the swap gate between two (d1)(d-1)-dimensional lattices. Then the path of QCA on a two-layer system
@@
interpolates between the identity and $v_{x}v_{x+1}^{\dagger}$, where $v_{x}$ acts as $v$ on the (d1)(d-1)-dimensional surface $\{x\}\times{\mathbb{Z}}^{d-1}$. Indeed, $u_{x,x+1}$ is a finite depth circuit, as we see by replacing the gates in the first $\mathrm{SWAP}$ by their conjugation by $v_{x}$. This operation preserves locality, so the result is still a finite depth circuit. Finally, the swindle $S(v,t)$ is the brickwork circuit of $u(v,t)$ gates shown in Fig. 3.
The swindle $S(v,t)$ demonstrates the pumping of the QCA $v$. On an length $L$ (even) system, the first layer of $u$ gates implements the QCA $v_{0}v_{1}^{\dagger}v_{2}v_{3}^{\dagger}\ldots v_{L}^{\dagger}$. The second layer of $u$ gates cancels all the $v$ factors in the bulk of the system, but leaves dangling unitaries $v_{0}v_{L}^{\dagger}$ acting on the left and right boundaries. The unitary $v$ has been pumped to the left of the system, leaving behind its inverse $v^{\dagger}$ on the right, thus proving 1.
The more precise statement capturing both 1 and 2 is that there is a homotopy equivalence between $\mathrm{QCA}_{d-1}$ and $\Omega\mathrm{QCA}_{d}$. That is, a map which—up to smooth deformations—has an inverse. We propose that the swindle $S$ is such an equivalence, and can sketch a proof as follows. Indeed, the homotopy inverse can be constructed through the pumping action of the loop $\gamma\in\Omega\mathrm{QCA}_{d}$. Restrict $\gamma$ to open boundaries (with a fixed choice of boundary for all $\gamma$), and define $\bar{S}(\gamma)=v$, where $v\in\mathrm{QCA}_{d-1}$ is the edge action induced by the restricted loop. It is straightforward to show that $\bar{S}$ can be chosen such that $\bar{S}S(v)=v$. Showing that $S\bar{S}$ can be continuously deformed to the identity is more involved, but we present an explicit deformation in Appendix B.
The more formal argument that $S$ is a homotopy equivalence for strictly local QCA is presented in Appendix B, assuming that any loop of QCA can be restricted to open boundary conditions. Making this argument rigorous requires verifying that the restriction to open boundaries always exists, which we leave for future work. (Our previous argument involving a Hamiltonian evolution is, rigorously, unsatisfying because it allows for exponential tails in the QCA to develop parallel to the boundary for $d>1$. Physically, this is not a problem, but it is an obstacle for proofs.)
III.3 Computing the classification
The $\Omega$-spectrum condition [Eq. (9)] implies that the bulk-boundary correspondence defined by pumping completely captures all multi-tone phases. In particular, iteratively applying Eq. (9), the classification of driven localized phases in any dimension is reduced to either the classification of static QCA, or a driven classification of zero-dimensional QCA,
@@ III.3 Computing the classification
The connected components $\pi_{0}(\mathrm{QCA}_{d})$ are known (for spin systems) for $d=1$ [58] and $d=2$ [61]. There are sound conjectures for $d=3$ [56, 109, 55]. The space $\mathrm{QCA}_{0}$ is a stabilization of the projective unitary group by stacking ancillae. We recall existing classifications and compute the homotopy groups of $\mathrm{QCA}_{0}$ in Sec. III.3.1 and Sec. III.3.2 respectively.
@@
Thus, we have all the homotopy groups $\pi_{n}(\mathrm{QCA}_{d})$ for spin systems with $d-n\leq 2$, and can make a strong conjecture for dn=3d-n=3. The results are listed in Tab. 1, and the corresponding classification of localized phases (excluding permutation QCA) is listed in Tab. 2. Comparing to the classification of invertible states shows that all the nontrivial phases in the tables are ALT phases.
$q$ 33 22 11 0 1-1 2-2 3-3 4-4 5-5
$h_{\mathrm{QCA}}^{q}$ Witt 0 $\log{\mathbb{Q}}^{\times}$ 0 $\mathbb{Q}$$\mathbb{Z}$} 0 $\mathbb{Q}$} 0 $\mathbb{Q}$}
$h_{\mathrm{inv}}^{q}$ 0 $\mathbb{Z}$} 0 0 0 $\mathbb{Z}$} 0 0 0
@@
Table 1: The homotopy groups $\pi_{n}(\mathrm{QCA}_{d})=:h_{\mathrm{QCA}}^{q}$ are a function of q=dnq=d-n, as is $\pi_{n}(\mathrm{inv}_{d})=:h_{\mathrm{inv}}^{q}$. For spin systems $h_{\mathrm{QCA}}^{3}$ is conjectured to be the Witt group of modular tensor categories [56, 109, 55, 110, 57]; $h_{\mathrm{QCA}}^{2}$ is trivial [61]; $h_{\mathrm{QCA}}^{1}$ is isomorphic to $\log{\mathbb{Q}}^{\times}$ [58]; and $h_{\mathrm{QCA}}^{q}$ for $q\leq 0$ is trivial for even $q$, $\mathbb{Q}$$\mathbb{Z}$} for q=1q=-1, and $\mathbb{Q}$} for odd $q\leq-3$. In all groups in the table, the group operation is addition.
dd\;\; n\;\;n 0 11 22 33 44 55
0 0 0 0 $\mathbb{Q}$} 0 $\mathbb{Q}$}
11 0 0 $\mathbb{Q}$$\mathbb{Z}$} 0 $\mathbb{Q}$} 0
22 0 $\log{\mathbb{Q}}^{\times}$ 0 $\mathbb{Q}$$\mathbb{Z}$} 0 $\mathbb{Q}$}
33 Witt 0 $\log{\mathbb{Q}}^{\times}$ 0 $\mathbb{Q}$$\mathbb{Z}$} 0
@@
Table 2: The classification of $n$-tone driven $d$-dimensional localized phases of spins with short-range entanglement. The classifying group differs from $\pi_{n}(\mathrm{QCA}_{d})$ for $(d,n)\in\{(0,1),(1,0)\}$ due to the non-uniqueness of the parameterization by QCA (Appendix A.3). All nontrivial elements of this table correspond to ALT phases.
III.3.1 Connected components
In general, the connected components $\pi_{0}(\mathrm{QCA}_{d})$ are not known. However, these groups have been calculated for $d\in\{1,2\}$ (and $d=0$, which is trivial), and the groups are known to be nontrivial in some higher dimensions. Note that while $\pi_{0}(X)$ has no canonical group structure for general spaces $X$, $\pi_{0}(\mathrm{QCA}_{d})$ inherits a group structure from either composition of QCA or tensor product. The resulting group operations are the same.
For one-dimensional QCA, the complete classification of $\pi_{0}(\mathrm{QCA}_{1})$ was obtained by Ref. [58], which also explicitly defined an invariant which diagnoses which connected component a given QCA belongs to. All one-dimensional QCA are path connected to a shift—a QCA which acts as a translation on some subset of the degrees of freedom. For example, in a chain of qubits, each QCA is connected to a QCA which translates all qubits by $r$ sites either right or left. For qubit chains, QCA are classified by the group $\mathbb{Z}$}.
However, we stabilize QCA by allowing ancillae of any Hilbert space dimension to be added to any site by tensor product. Thus, we must assume that the chain also has qutrit (three-state) degrees of freedom, and other higher level qudits. While a 4-level qudit can be decomposed into two qubits, and so does not give new invariants compared to the qubit system, qutrits and other prime qudits all produce new invariants. A shift of qutrits cannot be connected to a shift of qubits.
The resulting classifying group is succinctly presented as
$\pi_{0}(\mathrm{QCA}_{1})\cong\log{\mathbb{Q}}^{\times},$ (12)
[…]
$\pi_{0}(\mathrm{QCA}_{2})\cong 0.$ (14)
There is some subtlety in this statement. For QCA on a general two-dimensional manifold—that is, the lattice is obtained from a triangulation of a manifold—the classification may be nontrivial. However, all invariants relate to one-dimensional cycles in the manifold, and so belong to a lower dimensional part of the classification. For appropriate choices of topology [^2] This includes the topology used in Appendix B. on $\mathrm{QCA}_{2}$ (where the manifold is ${\mathbb{R}}^{2}$) this allows all two-dimensional QCA on the plane to be connected by continuous paths.
The classification of three dimensional QCA is still open. It is conjectured that $\pi_{0}(\mathrm{QCA}_{3})$ is isomorphic to the Witt group of modular tensor categories [56, 109, 55, 110]—that is, it has the same group structure as stacking two-dimensional topological orders, modulo an equivalence relation which identifies phases which may share a gapped edge. Presently, Clifford QCA (which map Pauli operators to Pauli operators) have been classified [109], and the classification is consistent with the Witt group conjecture for $\pi_{0}(\mathrm{QCA}_{3})$, but it has not been formally verified that all these phases of Clifford QCA remain nontrivial as non-Clifford QCA. Nonetheless, the existence of nontrivial elements of $\pi_{0}(\mathrm{QCA}_{3})$ is strongly supported by the constructions of Refs. [56, 55]. The models of those references can, indeed, be identified with two-dimensional topologically ordered phases. In fact, the MBL Hamiltonians corresponding to those models exhibit that topological order (specifically, the associated anyonic excitations) on their two dimensional surfaces [48, 56, 55, 110].
III.3.2 Zero dimensions
$\mathrm{QCA}_{0}(N)\cong PU(N).$ (15)
Unstable invariants.— We first recall the homotopy structure of projective unitaries for fixed $N$. The homotopy groups of $PU(N)$ for large but finite $N$ are all known,
There are even integral formulae for winding number invariants indexing the elements of these groups. These formulae involve first making a lift from projective unitaries to linear unitaries. Indeed, given any continuous map $V:S^{n}\to PU(N)$ (where $S^{n}$ is the $n$-dimensional sphere) with $n\neq 2$ it is possible to find a lift $\tilde{V}:S^{n}\to U(N)$ to the usual unitary group. (For $n=2$, we already have that $\pi_{2}(PU(N))\cong 0$, so we don’t need any formula to compute the homotopy class.) That is, if $q:U(N)\to PU(N)$ is the quotient map from the unitary group to the projective unitaries, we have $q\tilde{V}=V$. [^3] In general, the existence of such a lift is guaranteed by $V^{*}(\omega)\in H^{2}[S^{n},{\mathbb{Z}}]$ being trivial, where $\omega\in H^{2}[PU(N),{\mathbb{Z}}]$ is the Euler class for $q:U(N)\to PU(N)$ and $V^{*}$ is the map induced by $V$ on singular cohomology. But $H^{2}[S^{n},{\mathbb{Z}}]\cong 0$ for $n\neq 2$, so the map always exists for our cases of interest. Defining $A=(\mathrm{d}\tilde{V})\tilde{V}^{\dagger}$, where $\mathrm{d}$ is the exterior derivative on $S^{n}$, the winding number invariants are given by [31, 111]
$W_{n}[\tilde{V}]=C_{n}\int_{S^{n}}\mathrm{Tr}\left[A^{\wedge n}\right],$ (17)
$w_{1}[V]=W_{1}[\tilde{V}]\bmod N\in{\mathbb{Z}}_{N}$ (18)
Stable invariants.— Moving to larger $N$, the complication is that these winding numbers are invariant under the direct sum with an ancilla, not the tensor product. Indeed, denoting $\mathbbm{1}_{M}$ as the identity in $PU(M)$, we have from Eq. (17)
$W_{n}[V\otimes\mathbbm{1}_{M}]=MW_{n}[V].$ (19)
More generally, for $Q\in PU(M)$,
$W_{n}[V\otimes Q]=MW_{n}[V]+NW_{n}[Q].$ (20)
$\widetilde{W}_{n}[V]=C_{n}\int_{S^{n}}\frac{\mathrm{Tr}\left[A^{\wedge n}\right]}{\mathrm{Tr}\left[\mathbbm{1}\right]}.$ (21)
These are no longer integer invariants, but rather rational invariants (modulo 1 for $n=1$). Then Eq. (20) shows that $\widetilde{W}_{n}[V\otimes Q]=\widetilde{W}_{n}[V]+\widetilde{W}_{n}[Q]$.
Classification.— At a more formal level, for all $N,M\in{\mathbb{N}}$ we have injective homomorphisms
$\iota_{N\to NM}:PU(N)$ $\to PU(NM)$ (22)
$V$ $\mapsto V\otimes\mathbbm{1}_{M}$
[…]
with respect to this system. (We require that the multiplicity $n_{p}$ of every prime $p$ in the factorization $N=2^{n_{2}}3^{n_{3}}\ldots p^{n_{p}}\ldots$ approaches infinity, $n_{p}\to\infty$.) The homotopy groups of $\mathrm{QCA}_{0}$ can then be computed from the induced system
$\iota^{*}_{N\to NM}:\pi_{n}(PU(N))$ $\to\pi_{n}(PU(NM))$ (24)
$W_{n}[V]$ $\mapsto MW_{n}[V].$
By rescaling the winding numbers, we find that
[…]
Limited ancillae.— We obtained a classification by rationals because we allowed ancillae of any dimension to be appended to the system. Thus, all Hilbert space dimensions can occur in the denominator of $\widetilde{W}_{n}$. If, instead, we considered only a limited collection of Hilbert space dimensions for the ancillae, we would obtain a different classification. In fact, this classification will be a useful warm-up for the case of QCA with symmetry.
Say our available ancillae dimensions are given by
$\mathcal{R}=\{\rho\in{\mathbb{N}}\,:\,\text{ancilla of dimension }\rho\text{ exists}\},$ (30)
IV Symmetry enrichment
Physical systems frequently express symmetries which are crucial to their phenomenology, such as $U(1)$-symmetry leading to charge conservation. It is well known in many contexts that the presence of symmetry can affect topological classifications. Hamiltonians that were trivial may become nontrivial if some symmetry is imposed (called symmetry protection), or more generally phases without symmetry may fracture into several distinct phases if symmetry is introduced (called symmetry enrichment[72, 46, 16].